OpenStudy (anonymous):

If f(x)= 2x 2 + 3, then f(x - 1)= 2x 2 + 4x + 5. true or false?

OpenStudy (amoodarya):

f(x)= 2x 2 + 3 f(x-1)= 2(x-1)^ 2 + 3 = 2(x^2 -2x +1)+3 =2x^2 -4x +2+3 =2x^2-4x+5 not f (x - 1)= 2x 2 + 4x + 5